OpenStudy (anonymous):

Find the missing denominator. \frac{16}{32} = \frac{8}{?} 1 2 8 16

OpenStudy (anonymous):

i think its 16

OpenStudy (anonymous):

do 32 divided by 2

OpenStudy (anonymous):

u right

OpenStudy (anonymous):

thanks any thing els

OpenStudy (anonymous):

What is \frac{12}{16}in simplest form? \frac{4}{8} \frac{12}{16} \frac{3}{4} \frac{6}{8}

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

hurry up comp

OpenStudy (compassionate):

\( \frac{12}{16}\) What number can both of these be divided by? @Julianne6th , giving out answers is a quick way to get suspended from OpenStudy.

OpenStudy (anonymous):

you divide the top and botom by 2 untill you cant any more comp is wrong

OpenStudy (anonymous):

who cares about that @Compassionate

OpenStudy (compassionate):

@Minecraftdude1524 , if you're phishing for answers, you'll be suspended, too.

OpenStudy (anonymous):

i can just make a new account

OpenStudy (anonymous):

yeh @Compassionate

OpenStudy (anonymous):

i have 3 OS accounts

OpenStudy (anonymous):

@Minecraftdude1524 need anything els

OpenStudy (anonymous):

nope @Julianne6th

OpenStudy (anonymous):

ok well if you do just tag me

OpenStudy (anonymous):

k

OpenStudy (anonymous):

im a learner